OpenStudy (jenniferjuice):

Which equation represents the line passing through the points (3, 2) and (−9, 6)?

OpenStudy (anonymous):

first find the slope of the line using the formula: \[m=\frac{ y_{2}-y _{1} }{ x_{2}-x _{1} }\] then find the equation using the equation of the line:\[y-y _{1}=m(x-x _{1})\] where (x1, y1) is any of the given points..

OpenStudy (jenniferjuice):

4/-12

OpenStudy (jenniferjuice):

so ... y-2 = 4/-12 (x-3) ...?

OpenStudy (anonymous):

you can simplify more 4/-12..

OpenStudy (jenniferjuice):

y-2 = -1/3 (x-3)

OpenStudy (anonymous):

right! then simplify: y-2 = -1/3 (x-3) in terms of y which means, multiply -1/3 by the terms inside the bracket and transpose 2 to the right side and simplify.

OpenStudy (jenniferjuice):

wait im confused what do i do first?

OpenStudy (anonymous):